OpenStudy (anonymous):

Cylinder help..if the volume of a cylinder is 340 how do i find the perimeter, surface aria, and the hight

OpenStudy (anonymous):

You know that the Surface Area equation of a cylinder right? which is:\[SA = 2\pi r^2 + 2\pi r H\] and the volume is : \[V = \pi r^2 H\] Perimeter is : \[P = 2\pi R\] all you have to do is find a relationship between all of them and you'll be able to figure out the height, perimeter and SA , so you'll get : Notice that the volume is = pi r ^2 h and the perimeter is = 2pi r so the SA can be written as : \[SA = 2(340)H + PH\]

OpenStudy (anonymous):

are you sure that that's the only given?

OpenStudy (anonymous):

yep i am givin x 1-a2 i think that is the radious and i have to find all of those for numbers 1-12

OpenStudy (anonymous):

x = 1- a2?

OpenStudy (anonymous):

you need more given, I doubt that the only given is the volume lol

OpenStudy (anonymous):

ok so the whole thing is a project we have to figure out what is best for making a box..we are given x values 1-12 and the volume of 340 cubic inches 1-12 i believe are the radices

OpenStudy (anonymous):

i need to find the hight. the surface aria..and the perimeter of a cylinder

OpenStudy (anonymous):

with that info

OpenStudy (anonymous):

Alright, pick one of the radiuses and plug it in the Volume's equation to find H :) see if the number makes sense.

OpenStudy (anonymous):

ok so i take piXr squared and get the hight?

OpenStudy (anonymous):

for example .3.14X1 squared?

OpenStudy (anonymous):

V = pi r^2 H , you can plug in any value or R, then find H, and after finding H you can plug it in the perimeter, and when you find the perimeter you'll be able to find the SA, the new equation that I have wrote up there ^_^

OpenStudy (anonymous):

as far as I can go lol, since I don't have enough given

OpenStudy (anonymous):

im not getting how you find h when you need it in the equation

OpenStudy (anonymous):

alright, we have the volume right? :)

OpenStudy (anonymous):

\[V = 340 \]

OpenStudy (anonymous):

you have values of x, which are considered as the radius right?

OpenStudy (anonymous):

so you can take R = 4 for example and plug it in the following equation:\[V = \pi r^2h\] so we have R and V, we can solve for H, right?

OpenStudy (anonymous):

lol, are you getting the idea now? :)

OpenStudy (anonymous):

\[340=\pi(4) squared ?\]

OpenStudy (anonymous):

actually it's :\[H = (340)/(8\pi)\] and calculate it, that ofcourse if you choose your radius to be 4 ^_^ clear?

OpenStudy (anonymous):

wait it's 32 pi instead of 8 pi I forgot to square it lol!

OpenStudy (anonymous):

ok so i take x=340/r is this correct...if not i give up for the night and am going to bed..i am sure i am making this 100X harder than it actually is

OpenStudy (anonymous):

and that will give me the hight ?

OpenStudy (anonymous):

don't give up lol, alright, let's try once more! WE have x values, right?

OpenStudy (anonymous):

yes those values= the radius...so if the virst x value is 1 how do i get the hight?

OpenStudy (anonymous):

it's simple, just take your time and concentrate with me

OpenStudy (anonymous):

Okay so x = r = 1 we have given: R = 1 V = 340 find H= ?

OpenStudy (anonymous):

yep thats what i need to do...now how?

OpenStudy (anonymous):

The equation is : \[V = \pi r^2 h\]

OpenStudy (anonymous):

plug in the given in the equation and solve for H : 340 = pi (1)^2 H H = 340/pi H = 108.23 in ^_^

OpenStudy (anonymous):

we have found the height now :)

OpenStudy (anonymous):

now we need to find the perimeter right?

OpenStudy (anonymous):

are you following me lol? =P wakey wakey~

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

okay ^_^

OpenStudy (anonymous):

ok so the hight for a radios of 1 is 108.2 correct

OpenStudy (anonymous):

excellent!

OpenStudy (anonymous):

ok so now perimeter

OpenStudy (anonymous):

now, we can find the perimeter right? The equation is : P = 2pi R we have R = 1, solve :)

OpenStudy (anonymous):

ok so i take \[2piX(1)squared\]

OpenStudy (anonymous):

i got 6.28

OpenStudy (anonymous):

lol no, the equation is :\[P = 2 \pi R\] substitute R in the equation and you'll get P ^_^

OpenStudy (anonymous):

lol yeah , that's right :)

OpenStudy (anonymous):

now, we can find the SA right?

OpenStudy (anonymous):

ok so now the SA

OpenStudy (anonymous):

The equation is :\[SA = 2\pi r^2 + 2\pi r h\] we have R = 1 and H = 108.2 substitute the given and solve ^_^

OpenStudy (anonymous):

i got 686.1

OpenStudy (anonymous):

very good !

OpenStudy (anonymous):

you have solved the question using R = 1 ^_^ did you get the idea now? :)

OpenStudy (anonymous):

ok let me make sure i can get this for 2 now and then ill let you go..if you dont mind

OpenStudy (anonymous):

good morning

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

alright, it's the same idea blackford, but different values for R ^_^, good morning andy

OpenStudy (anonymous):

same way of solving , good luck! ^_^

OpenStudy (anonymous):

i got 120.8

OpenStudy (anonymous):

for?

OpenStudy (anonymous):

the hight with a radios of 2

OpenStudy (anonymous):

hmm it should be 27.06 (340)/(4pi) = H H = 27.06 in

OpenStudy (anonymous):

and andy going quite lol

OpenStudy (anonymous):

oops i see i think..i just realized you want h by itself

OpenStudy (anonymous):

yep ^_^ since we want to find it

OpenStudy (anonymous):

ok for a R of 3 i got 12.02 for the hight..correct?

OpenStudy (anonymous):

excellent ^_^

OpenStudy (anonymous):

ok i think i am getting it now

OpenStudy (anonymous):

you got the idea now :)

OpenStudy (anonymous):

you can proceed on your own right?

OpenStudy (anonymous):

for now thank you :)

OpenStudy (anonymous):

np ^_^

OpenStudy (anonymous):

i got a new one

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

now for a square pyramid. this time i need the hight,perimeter, SA, and lateral aria

OpenStudy (anonymous):

volume still 340 and same x values 1-12

OpenStudy (anonymous):

All you have to do is write down the new equations for the square pyramid

OpenStudy (anonymous):

i think that x= base length now

OpenStudy (anonymous):

Volume is : \[V = (b^2h)/3\] Surface Area is : \[SA = 2bs + b^2\] and perimeter : \[P = (\S-B)(1/2)(l)\] find the relationship and solve ^_^ same idea as the previous one :)

OpenStudy (anonymous):

P = (S-B)(1/2)(L) *

OpenStudy (anonymous):

oh and the lateral area is = PL/2 where P = perimeter and L = slant height ^_^

OpenStudy (anonymous):

ok so 340=1squared(h)/3

OpenStudy (anonymous):

oh yeah i want the slant hight not lateral aria

OpenStudy (anonymous):

so \[340=1^2(h).....340=1/3 =.33 340/.33..correct or no\]

OpenStudy (anonymous):

340(3) = h h = 1020

OpenStudy (anonymous):

lol, I see what's your problem, you got the idea, but have trouble with calculation, take your time answering the questions :) step at a time ^_^